OpenStudy (anonymous):

(9-sqrt{2})(6+sqrt{2}) is the answer 15sqrt{2}?

OpenStudy (anonymous):

you gotta multipy this out

OpenStudy (anonymous):

\[(9-\sqrt{2})(6+\sqrt{2})\] do i use the foil method?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i dont think this is right but here it is \[54+9\sqrt{2}-6\sqrt{2}-\sqrt{2}\]

OpenStudy (anonymous):

everything is right, excpect for the last part, its not sqrt(2), its just 2

OpenStudy (anonymous):

now do i collect the like terms and it will be\[54+\sqrt{2}-\sqrt{2}\] so will the answer be 54?

OpenStudy (anonymous):

56?

OpenStudy (anonymous):

no, it would be 54+3sqrt(2)-2=56+3sqrt(2)

OpenStudy (anonymous):

ok so the final answer would be \[56+3\sqrt{2}\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

well i got 1% of it right lol

OpenStudy (anonymous):

you mean its wrong?

OpenStudy (anonymous):

oh i ment i just got the 56 right well no i got that wrong too so i didnt get much of it right

OpenStudy (anonymous):

sorry 52+3sqrt(2)

OpenStudy (anonymous):

come on you have to checking a well, dont just trust people on here, we make mistakes

OpenStudy (anonymous):

well thats why i ask to check my work cause when i do it i get different answers, and i dont ask to have the answers just told to me cause i still wont know how to work it,

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thats why i ask if the steps im doing is right

OpenStudy (anonymous):

but 9*6=54 how is it 52? did you subtract the 2